OpenStudy (eg145341):

The floor of a bedroom is 12 feet by 15 feet and the walls are 7 feet high. One gallon of paint covers 250 square feet. How many gallons of paint do you need to paint the walls and the ceiling of the bedroom? Round your answer to the whole gallon.

OpenStudy (radar):

We need to make some assumptions" Assumption 1. The ceiling is the same size as the floor (12 X 15) or 180 sq ft. Assumption 2. We are going to paint any doors or windows that exists in the wall. Are you O.K. with these assumptions?

OpenStudy (anonymous):

Ceiling: \[12 \times 15 = 180\]Walls:\[(12 \times 7 \times 2)+ (15 \times 7 \times 2)=378\]Total area: \[180+378=558\] Total gallons:\[\frac{558}{250}=2.232\]